in an isosceles triangle, the third side is 50% of the length of one of the equal sides. Find the length of each side when the perimeter is 125ft. i know the answer i just dont know the equation?
Answer by mananth(16949)

in an isosceles triangle, the third side is 50% of the length of one of the equal sides. Find the length of each side when the perimeter is 125ft. i know the answer i just dont know the equation?
let the length of the equal side be x
The other side will also be x
Thethird side = 1/2 x
x+x+x/2 = 125
2x+2x +x / 2 = 125
5x/2 = 125
5x= 250
x=50 feet are the two sides
the third side = 25